satellite73 (satellite73):

Multiply \[24\times 53\] (or any two 2-digit numbers) using 3 multiplications instead of the usual 4

OpenStudy (anonymous):

no takers on this one?

OpenStudy (saifoo.khan):

No Takers, Only Under Takers!!

OpenStudy (anonymous):

i think this is a computer algorithm problem. cuts down on multiplications

OpenStudy (asnaseer):

I can do it using only two multiplications: 20 x 53 plus 4 times 53 oh - ok - misunderstood the question...

OpenStudy (anonymous):

yeah that is still 4 right?

OpenStudy (anonymous):

use ethiopian multiplication!

OpenStudy (anonymous):

ok, what might that be?

OpenStudy (anonymous):

rosettacode.org/wiki/Ethiopian_multiplication

OpenStudy (anonymous):

http://rosettacode.org/wiki/Ethiopian_multiplication

OpenStudy (anonymous):

wow! not exactly what i had in mind though...

OpenStudy (anonymous):

method i had was somewhat shorter

OpenStudy (asnaseer):

here is a list of algorithms for doing multiplication: http://en.wikipedia.org/wiki/Multiplication_algorithm

OpenStudy (anonymous):

funny i looked here and didn't see it

OpenStudy (asnaseer):

One of these is particularly interesting: http://en.wikipedia.org/wiki/Karatsuba_algorithm

OpenStudy (anonymous):

hint is \[(a+b)(c+d)=ac+ad+bc+bd\] and you need the numbers \[ac, bd, ad+bc\]

OpenStudy (anonymous):

yeah that is it! for example \[53\times 24\] \[5\times 2=10,3\times 4=12,(5+3)(2+4)=8\times 6=48, 48-10-12=26\] 1 2 26 10 __________-- 1 2 7 2

OpenStudy (asnaseer):

yes - until you brought this up I was wasn't even aware of efficient multiplication methods. thanks for opening the door to new knowledge. :-)

OpenStudy (anonymous):

i never knew the name, and only remember it from a discrete algorithms class. in fact it is all i recall from that class

OpenStudy (anonymous):

karatsuba! i will try to remember that name

OpenStudy (asnaseer):

how can we ever forget a name like that! :-)

OpenStudy (anonymous):

especially with a bookmark!

OpenStudy (asnaseer):

lol!

OpenStudy (asnaseer):

I have just invented a way for me to remember this: rabbits hop around and multiply very fast. so what is their favourite meal? "carrot soup" of course! (hence katatsuba)